Cost of Units Transferred Out and Ending Work in Process

The costs per equivalent unit of direct materials and conversion in the Rolling Department of Kraus Steel Company are $0.50 and $1.65, respectively. The equivalent units to be assigned costs are as follows:

Equivalent Units
Direct Materials Conversion
Inventory in process, October 1 0 3,200
Started and completed during October 46,000 46,000
Transferred out of Rolling (completed) 46,000 49,200
Inventory in process, October 31 5,000 2,500
Total units to be assigned costs 51,000 51,700

The beginning work in process inventory on October 1 had a cost of $1,860. Determine the cost of completed and transferred-out production, the ending work in process inventory, and the total costs assigned by the Rolling Department.

Completed and transferred-out production $
Inventory in process, October 31 $
Total costs assigned by the Rolling Department $
Answer #1

Solution

Completed and transferred-out production $ 106,040
Inventory in process, October 31 $ 6,625
Total costs assigned by the Rolling Department $ 112,665

Working

Cost per Equivalent Units AND Cost to be accounted for
Material Conversion cost TOTAL
Cost incurred in Current Period $ 25,500 $ 85,305 $ 110,805
Total Equivalent Units                                              51,000                                                       51,700
Cost per Equivalent Units $ 0.50 $ 1.65 $ 2.15
Cost of Beginning WIP $ 1,860
Total cost to be accounted for $ 112,665

.

Cost report [Including Reconciliation]
Cost of Units Transferred               49,200.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Units Transferred TOTAL
A From Beginning WIP:
Material                            -   $ 0.50 $ 0
Conversion cost                     3,200 $ 1.65 $ 5,280 $ 5,280
B From units started & completed
Material                  46,000 $ 0.50 $ 23,000
Conversion cost                  46,000 $ 1.65 $ 75,900 $ 98,900
C=A+B Total $ 104,180
D Cost of Beginning WIP $ 1,860
E=C+D Total Cost of Units transferred $ 106,040
Cost of Ending WIP                 5,000.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Ending WIP TOTAL
F Material                     5,000 $ 0.50 $ 2,500
G Conversion cost                     2,500 $ 1.65 $ 4,125
H = F + G Total Cost of Ending WIP $ 6,625
I = E + H Total Cost accounted for $ 112,665
J Rounding Off Differences $ 0
K = I + J Total cost to be accounted for [equals to computed in STEP 3] $ 112,665